2x + 6y = 50
10x + 0y = 10
Solve by the substitution method.

Answer by mananth(16946)

2x + 6y = 50
10x + 0y = 10
10x+0=10
10x=10
/10
x= 1
...
plug value of x in first equation
2x+6y= 50
2*1+6y=50
2+6y=50
-2 -2
6y=50-2
6y=48
/6
6y/6 = 48/6
y =8
m.ananth@hotmail.ca